Solve by any method
-4x + 3y = 2
5x - 2y = -3

Answer by checkley71(8403)

-4X+3Y=2 MULTIPLY THIS EQUATION BY 5 WE GET -20X+15Y=10
5X-2Y=-3 MULTIPLY THIS EQUATION BY 4 WE GET 20X-8Y=-12
NOW ADD THESE TWO EQUATIONS THUS
-20X+15Y=10
20X-8Y=-12
-------------------------------------
7Y=-2
Y=-2/7 THEN
-4X+3*-2/7=2
-4X-6/7=2
-4X=2+6/7
-4X=20/7
X=(20/7)/-4
X=(20/7)(-1/4)
X=-5/7
PROOF
5(-5/7)-2(-2/7)=-3
-25/7+4/7=-3
(-25+4)/7=-3
-21/7=-3
-3=-3